Using the alternate form of Newton's second law, compute the force (in N) needed to accelerate a 1,020 kg car from 0 to 15 m/s in 17 s. The change in momentum is the car's momentum when traveling 15 m/s minus its momentum when going 0 m/s. (Enter the magnitude.)
Answer #1

Newton's second law states that the force acting on a body is equal to the rate of change of momentum.

So, initial momentum = mass * initial velocity = 1020 * 0 = 0 kg-m/sec

Also, final momentum = mass * final velocity = 1020 * 15 = 15300 kg-m/sec

Now, according to second law:

On solving,

Force = 900 N

PLEASE RATE IF YOU LIKE IT.